• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Camsonne

    Log entry time 03:55:27 on August30,2009

    Entry number 286989

    keyword=Compton electron detector run 18475

    Rate is about 150 Hz with laser on and about 70 hertz with laser off.
    Threshold is 16 and the position is 4 mm, Compton edge around strip 18. It seems I miss half of the detector. Power supply seems OK, I will have to check the threshold / HV distribution board, I guess I will swap it at the next access. We see triple coincidence already.
    Quick efficiency analysis indicates efficiencies 24 % for plane B, 18.6 % for plan C and 13.9% for plan D. Plan A is disabled since it has only an analog board.

    FIGURE 1

    FIGURE 2